OpenStudy (jonny):

rearrange p=2a+c make a the subject

OpenStudy (jonny):

i have done p=2a+c p-c=2a Then what?

OpenStudy (shubhamsrg):

divide both sides by 2 and see the magic!! ;)

OpenStudy (jonny):

so its: a= 2/p-c

OpenStudy (shubhamsrg):

hmmn,,i wont comment much i'l tell you the ans its a= (p-c)/2

OpenStudy (jonny):

whys that the answer

OpenStudy (shubhamsrg):

you divide both sides by 2 you had 2a = p - c then we have (2a/2) = (p-c)/2 => a = (p-c)/2 was i clear enough?